OpenStudy (anonymous):

find the slope and y-intercept of each line: 3x-4y=12

OpenStudy (anonymous):

slope-yintercept form is mx+b = y, so: 3x-4y=12 3x-12 = 4y (3/4)x-3 = y slope = m = 3/4 y intercept = b = -3